OpenStudy (anonymous):

The probability of Dawn reading before going to bed is 5/6. The probability of Dawn brushing her teeth before going to bed is 6/7. The probability of Dawn watching TV before going to bed is 1/5. What is the probability that she will do all three activities before going to bed?

OpenStudy (amistre64):

does any one activity affect the probability of the other?

OpenStudy (anonymous):

Yes, i think

OpenStudy (amistre64):

i dont think they have any affect on each other, i believe that they are independant events. what in your material suggests a way that we can find P(A and B and C)

OpenStudy (anonymous):

Add them?

OpenStudy (amistre64):

almost ... adding is associated with the word 'or'. if the events are independant, and we want to 'and; them; we can simply multiply them togther.

OpenStudy (amistre64):

so, the way i see it, we can simply do this\[P(A\cap B\cap C)=\frac{5}{6}*\frac{6}{7}*\frac{1}{5}\]

OpenStudy (anonymous):

How would i multiply the fractions together? Would I cross multiply?

OpenStudy (amistre64):

no, fractions are just multiplied straight across. cross multiplication is a memory device used when there are fractions seperated by an = sign. none of our fractions are seperated by an = sign in this

OpenStudy (anonymous):

Okay. Soooo

OpenStudy (amistre64):

soo, thats as far as im willing to take you on this. your question was on probability, not mutliplication of fractions.

OpenStudy (anonymous):

It would be 1/7?

OpenStudy (amistre64):

it would be yes :)

OpenStudy (anonymous):

A bag contains 3 red marbles, 4 blue marbles, and 5 white marbles. What is the probability of pulling out a white marble, and then a blue marble if none are replaced?